# Learning to Route: A Rule-Driven Agent Framework for Hybrid-Source Retrieval-Augmented Generation

### v1

@@ Abstract.
Large Language Models (LLMs) have shown remarkable performance on general Question Answering (QA), yet they often struggle in domain-specific scenarios where accurate and up-to-date information is required. Retrieval-Augmented Generation (RAG) addresses this limitation by enriching LLMs with external knowledge, but existing systems primarily rely on unstructured documents, while largely overlooking relational databases, which provide precise, timely, and efficiently queryable factual information, serving as indispensable infrastructure in domains such as finance, healthcare, and scientific research. Motivated by this gap, we conduct a systematic analysis that reveals three central observations: (i) databases and documents offer complementary strengths across queries, (ii) naively combining both sources introduces noise and cost without consistent accuracy gains, and (iii) selecting the most suitable source for each query is crucial to balance effectiveness and efficiency. We further observe that query types show consistent regularities in their alignment with retrieval paths, suggesting that routing decisions can be effectively guided by systematic rules that capture these patterns. Building on these insights, we propose a rule-driven routing framework. A routing agent scores candidate augmentation paths based on explicit rules and selects the most suitable one; a rule-making expert agent refines the rules over time using QA feedback to maintain adaptability; and a path-level meta-cache reuses past routing decisions for semantically similar queries to reduce latency and cost. Experiments on three QA benchmarks demonstrate that our framework consistently outperforms static strategies and learned routing baselines, achieving higher accuracy while maintaining moderate computational cost.
Routing Mechanisms, Database, Retrieval-Augmented Generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Data management systemsccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
(a) Retrieval Path Effectiveness is Question-Dependent
(b) Excess Context Can Hurt Accuracy
(c) Balancing Accuracy and Efficiency
(d) Problem Types Align with Retrieval Paths
Figure 1. Problem Analysis
Large language models (LLMs) have achieved remarkable success across a wide range of natural language processing tasks, particularly in question answering (QA). Yet, despite their strong capabilities, LLMs still struggle in specialized QA scenarios such as enterprise knowledge access or domain-specific expert consultation, where accurate and up-to-date information is required but may not be captured in the model’s parametric memory. To address this limitation, Retrieval-Augmented Generation (RAG) has been proposed (Borgeaud et al., 2022; Izacard et al., 2023; Lewis et al., 2020), which enriches LLMs with external knowledge retrieval, thereby improving factual grounding and adaptability.
@@
To better understand the role of relational databases in RAG, we conduct a set of motivating experiments on the TATQA (Zhu et al., 2021) dataset using GPT-4.1-mini. Here, augmentation follows the standard RAG paradigm: the retrieved evidence is appended to the question and provided to the LLM for answer generation, with the only difference being the external knowledge source. i) Complementarity. We first compare two augmentation paths: one enhanced by relational databases (DB), where query results from relational tables are verbalized into text and fed to the LLM, and the other by unstructured documents (Doc), where relevant passages are retrieved from a document corpus. Figure 1(a) visualizes per-question outcomes, where each cell corresponds to one query and the color denotes which augmentation path yields the correct answer. The results reveal strong complementarity: many questions can only be answered correctly by DB augmentation, while others require Doc augmentation. Neither source dominates across all queries, and relying on a single source leaves substantial coverage gaps. ii) Naive Hybridization. A natural follow-up is to combine DB and Doc evidence simultaneously by concatenating both sources and passing them to the LLM. Figure 1(b) further shows per-question outcomes under this hybrid strategy, where the additional color indicates cases in which both sources are combined. We observe that many queries that can be correctly answered by a single path (DB or Doc) may fail under hybrid augmentation. Instead, redundant or noisy evidence often distracts the model, leading to incorrect answers (see Section 5.8 for details), while the token count grows dramatically, increasing inference latency and monetary cost. Thus, simply ”feeding more” is neither efficient nor reliable. iii) Necessity of Routing. Finally, we examine the impact of applying a uniform strategy across all queries. Figure 1(c) demonstrates that single-source inputs limit accuracy, while hybrid inputs incur high overhead. In contrast, an ideal system should allocate each query to its best-suited path, simultaneously improving accuracy and keeping token usage moderate. This motivates the need for a routing mechanism that achieves a better balance between effectiveness and efficiency.
@@
Beyond aggregate performance, we further observe that queries are not randomly distributed across sources but instead exhibit clear structural regularities. In particular, fact-centric or numerical questions are typically better served by database augmentation, whereas open-ended or descriptive queries align more naturally with document retrieval, a pattern consistently observed across multiple datasets (see Figure 1(d) for illustration). This observation suggests that routing can move beyond opaque or black-box classifiers and instead be grounded in explicit, interpretable rules that capture these understandable patterns. Such a rule-driven perspective also mirrors human decision-making. Cognitive studies show that people rarely exhaust all possible options; rather, they often apply simple ”if–else” heuristics to make efficient choices under resource constraints (Gigerenzer et al., 2000; Simon, 1990). For example, if it rains, people carry an umbrella; if the weather is clear, they leave it at home. These everyday heuristics clearly illustrate how humans rely on interpretable rules to balance accuracy and efficiency without overcomplicating decisions. Motivated by this analogy, we design our framework around rule-driven routing, where transparent rules encode observed query-path alignments and are incrementally refined through feedback to balance accuracy, efficiency, and interpretability.
Our framework incorporates three key components: (i) a Rule-Driven Routing Agent, which avoids the inefficiency of static augmentation by evaluating candidate paths with interpretable rules and selecting the most suitable one; (ii) a Rule-Making Expert Agent, which refines these rules using feedback from QA performance, thereby reducing reliance on static, hand-crafted rules and enabling adaptation to different datasets; and (iii) a Path-Level Meta-Cache, which further accelerates inference by reusing routing decisions for repeated or semantically similar queries, eliminating unnecessary agent calls. Together, these components enable adaptive, interpretable, and efficient integration of structured and unstructured knowledge sources.
Our contributions are summarized as follows: i) We conduct analysis of routing between relational databases and document corpora in RAG, highlighting their complementary strengths. ii) We introduce a rule-driven routing framework that integrates a routing agent, a rule-making expert agent, and a path-level meta-cache, aiming to balance accuracy, interpretability, and efficiency. iii) We evaluate the proposed framework on three QA benchmarks, showing that it achieves consistent improvements over static strategies and alternative routing baselines, while maintaining moderate cost.
@@ 2.1. Hybrid Knowledge Sources for RAG
Retrieval-Augmented Generation (RAG) is an influential framework that enhances the capabilities of large language models (LLMs) by enabling access to external knowledge sources, thereby improving their performance on tasks that require domain-specific information (Borgeaud et al., 2022; Izacard et al., 2023; Lewis et al., 2020). Unlike traditional language models, which generate text based solely on internal parameters, RAG models incorporate a retrieval component that fetches relevant content from external data sources and integrates it into the generation process. Traditional RAG systems primarily utilize large-scale unstructured corpora, such as Wikipedia and web documents, as their external knowledge base (Karpukhin et al., 2020; Ram et al., 2023; Yao et al., 2023). Building on this foundation, more recent work has expanded the spectrum of retrieval sources to further enhance RAG’s capabilities. Some studies integrate knowledge graphs, which provide structured representations of entities and relationships. Knowledge graphs enable more context-aware and entity-centric retrieval, supporting complex information needs and facilitating multi-hop reasoning (Ma et al., 2025; Sarmah et al., 2024; Li et al., 2024). Other approaches incorporate multi-modal data—such as images, videos, and tables—enabling RAG systems to synthesize information across modalities and improving their robustness on heterogeneous, complex queries (Abootorabi et al., 2025; Hu et al., 2023; Chen et al., 2022). In contrast, relational databases store highly structured and precise factual data, supporting efficient updates and fast queries (Harrington, 2016; Codd, 2007). Despite their widespread use in many enterprise and scientific applications, the integration of relational databases as knowledge sources in RAG has received relatively little attention (Seabra et al., 2024). There is a lack of systematic analysis and methodical approaches for combining the complementary strengths of relational databases and unstructured documents within RAG frameworks. Addressing this gap, our work aims to explore and propose novel strategies for the joint utilization of these two distinct types of data sources in RAG.
@@ 2.2. Routing Mechanisms in RAG
Routing mechanisms have received increasing attention in LLMs (Wang et al., 2025) and RAG, enabling dynamic and intelligent decision-making about how each query is processed. Unlike fixed pipelines, routing-enhanced RAG architectures employ dedicated modules that analyze the characteristics of incoming queries and adaptively select the most appropriate retrieval and reasoning pathways, thereby improving both efficiency and response quality. A fundamental role of routing in RAG is to determine the appropriate complexity of the retrieval pipeline for each query. While some questions can be resolved through simple, single-hop retrieval, others require more complex, multi-hop reasoning or iterative retrieval procedures (Jeong et al., 2024; Tan et al., 2024; Maekawa et al., 2024; Yin et al., 2025). Adaptive routing allows the system to dynamically decide whether external retrieval is needed or if the language model’s parametric knowledge suffices, effectively balancing computational cost and answer accuracy. Another crucial dimension of routing concerns the choice of retrieval methods. Early RAG systems often used simple rule-based or classifier-driven approaches, sending queries with strong lexical overlap to sparse retrievers such as BM25 (Robertson et al., 1994), while directing semantically complex or ambiguous queries to dense retrievers (Karpukhin et al., 2020). More recent advances have introduced lightweight classifiers, neural models, and multi-agent frameworks that can select or combine multiple retrieval strategies, sometimes explicitly modeling the interplay between external retrievals and the language model’s internal knowledge (Şakar and Emekci, 2025; Zhang et al., 2025). Beyond retrieval strategy, routing also governs data source selection. In real-world scenarios, the knowledge necessary to answer a query may reside in document collections, knowledge bases, or other sources. Modern RAG systems increasingly deploy specialized routers—such as domain routers or dynamism routers—to intelligently map queries to the most relevant data sources and to orchestrate hybrid retrieval workflows that capitalize on the complementary advantages of each data type (Wu et al., 2025; Guerraoui et al., 2025; Xia et al., 2025). In this work, we focus on dynamic routing and the integration of structured relational databases with unstructured documents in RAG systems. By recognizing and leveraging the distinct strengths of these sources, we propose a routing mechanism that adaptively selects the knowledge source.
@@ 3. Preliminaries
Database-Enhanced QA. We further consider advanced database-enhanced QA, which leverages structured relational data to support highly time-sensitive and factually reliable answer generation. For a given question $q$, the system first retrieves relevant tables by matching $q$ against a carefully pre-built index of table metadata (including schemas, high-frequency values, and LLM-generated descriptions). After candidate tables are successfully identified, a structured query (e.g., SQL) is automatically generated and executed to extract a precise set of factual records $\mathcal{F}_{q}$ from the database. These records, as immutable evidence, are then incorporated into the prompt: $a=\mathrm{LLM}(q,\mathcal{F}_{q})$. In our implementation, we explicitly instruct the LLM that these retrieved database facts are not to be altered, thereby ensuring high precision, reliability, and full auditability of the final answer.
@@
Routing Mechanisms. The complementary nature of document and database augmentation presents a new challenge: for each incoming question, how to dynamically select the most effective enhancement pathway, so as to balance answer quality, response latency, and computational cost. We formalize this as a query routing problem: given a question $q$, a routing agent selects an augmentation path $p_{q}\in\mathcal{P}=\{\texttt{Doc},\texttt{DB},\texttt{Hybrid},\texttt{LLM}\}$ according to a policy, which aims to optimize an overall utility objective combining quality, latency, and cost. The selected pathway then determines which information is retrieved and used to construct the prompt for the downstream LLM, which generates the final answer. This design allows the system to flexibly adapt to each query’s needs, leveraging the complementary strengths of both data sources while managing system efficiency.
@@ 4.2. Rule-Driven Routing Agent
Expert-Initialized Rules. Our routing design begins with a set of interpretable rules $\mathcal{R}$, initialized by domain experts. Each rule assigns additive scores to candidate augmentation paths based on recognizable query patterns. Formally, we denote by $S_{p}(q)$ the score assigned to path $p$ for query $q$, where higher values indicate stronger preference for selecting that path. A subset of illustrative rules (not exhaustive) is as follows:
@@ 4.3. Rule-Making Expert Agent
4.4. Path-Level Meta-Cache
While our rule-driven routing agent substantially reduces the inefficiency of naive hybrid augmentation, it still requires an additional agent call for rule evaluation. This extra step introduces some overhead, yet the cost is minor compared to the substantial savings achieved by avoiding redundant hybrid augmentation. Nevertheless, in large-scale deployments even small latencies can accumulate, motivating further optimization. Since many real-world queries are repeated or semantically similar, caching becomes a natural choice to further reduce overhead. However, in dynamic scenarios, such as those involving frequently updated relational databases, traditional answer-level caches may become unreliable. For example, caching answers to queries like ”What is the revenue growth over the past 7 days?” can easily lead to stale or incorrect results as the underlying data changes. To overcome this limitation, we propose a path-level meta-cache that operates at the routing decision level rather than the answer level.
(5) $\text{sim}(z_{q_{j}},z_{q^{\prime}})\geq\tau,$
5. Experiment

### v2

@@ Abstract.
Large Language Models (LLMs) have shown remarkable performance on general Question Answering (QA), yet they often struggle in domain-specific scenarios where accurate and up-to-date information is required. Retrieval-Augmented Generation (RAG) addresses this limitation by enriching LLMs with external knowledge, but existing systems primarily rely on unstructured documents, while largely overlooking relational databases, which provide precise, timely, and efficiently queryable factual information, serving as indispensable infrastructure in domains such as finance, healthcare, and scientific research. Motivated by this gap, we conduct a systematic analysis that reveals three central observations: (i) databases and documents offer complementary strengths across queries, (ii) naively combining both sources introduces noise and cost without consistent accuracy gains, and (iii) selecting the most suitable source for each query is crucial to balance effectiveness and efficiency. We further observe that query types show consistent regularities in their alignment with retrieval paths, suggesting that routing decisions can be effectively guided by systematic rules that capture these patterns. Building on these insights, we propose a rule-driven routing framework designed specifically for hybrid-source RAG. A routing agent scores candidate augmentation paths based on explicit rules and selects the most suitable one; a rule-making expert agent refines the rules using QA feedback to produce more comprehensive and reliable decision criteria; and a path-level meta-cache reuses past routing decisions for semantically similar queries to reduce latency and cost. Experiments on three QA datasets demonstrate that our framework consistently outperforms static strategies and learned routing baselines, achieving higher accuracy while maintaining moderate computational cost.
Routing Mechanisms, Database, RAG
copyright: noneconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesccs: Information systems Data management systemsccs: Information systemsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
(a) Complementarity
(b) Naive Hybridization
(c) Advantages of Routing
(d) Pattern Analysis
Figure 1. Problem Analysis. (a) Retrieval Path Effectiveness is Question-Dependent. (b) Excess Context Can Hurt Accuracy. (c) Routing Makes Balance Accuracy and Efficiency. (d) Problem types align with database- vs. document-oriented retrieval.
Large language models (LLMs) have achieved remarkable success across a wide range of natural language processing tasks (Shao et al., 2026; Zhai et al., 2026), particularly in question answering (QA) (Liu et al., 2025). Yet, despite their strong capabilities, LLMs still struggle in specialized QA scenarios such as enterprise knowledge access or domain-specific expert consultation, where accurate and up-to-date information is required but may not be captured in the model’s parametric memory. To address this limitation, Retrieval-Augmented Generation (RAG) has been proposed (Borgeaud et al., 2022; Izacard et al., 2023; Lewis et al., 2020), which enriches LLMs with external knowledge retrieval, thereby improving factual grounding and adaptability.
@@
To better understand the role of relational databases in RAG, we conduct a set of motivating experiments on the TATQA (Zhu et al., 2021) dataset using GPT-4.1-mini. Here, augmentation follows the standard RAG paradigm: the retrieved evidence is appended to the question and provided to the LLM for answer generation, with the only difference being the external knowledge source. i) Complementarity. We first compare two augmentation paths: one enhanced by relational databases (DB), where query results from relational tables are verbalized into text and fed to the LLM, and the other by unstructured documents (Doc), where relevant passages are retrieved from a document corpus. Figure 1(a) visualizes per-question outcomes, where each cell corresponds to one query and the color denotes which augmentation path yields the correct answer. The results reveal strong complementarity: many questions can only be answered correctly by DB augmentation, while others require Doc augmentation. Neither source dominates across all queries, and relying on a single source leaves substantial coverage gaps. ii) Naive Hybridization. A natural follow-up is to combine DB and Doc evidence simultaneously by concatenating both sources and passing them to the LLM. Figure 1(b) further shows per-question outcomes under this hybrid strategy, where the additional color indicates cases in which both sources are combined. We observe that many queries that can be correctly answered by a single path (DB or Doc) may fail under hybrid augmentation. Instead, redundant or noisy evidence often distracts the model, leading to incorrect answers, while the token count grows dramatically, increasing inference latency and monetary cost. Thus, simply ”feeding more” is neither efficient nor reliable. iii) Necessity of Routing. Finally, we examine the impact of applying a uniform strategy across all queries. Figure 1(c) demonstrates that single-source inputs limit accuracy, while hybrid inputs incur high overhead. In contrast, an ideal system should allocate each query to its best-suited path, simultaneously improving accuracy and keeping token usage moderate. This motivates the need for a routing mechanism that achieves a better balance between effectiveness and efficiency (Wang et al., 2025; Shao et al., 2024).
@@
Beyond aggregate performance, we further observe that queries are not randomly distributed across sources but instead exhibit clear structural regularities. In particular, fact-centric or numerical questions are typically better served by database augmentation, whereas open-ended or descriptive queries align more naturally with document retrieval, a pattern consistently observed across multiple datasets (see Figure 1(d) for illustration). However, existing learned routers—whether classifier-based or LLM-based—often struggle to stably capture these heterogeneous patterns. They require large labeled data to train, behave as black boxes during deployment, and tend to produce uncontrollable routing. These limitations motivate a more transparent and rule-grounded approach, where routing decisions are guided by explicit and interpretable rules that encode the observed query–path regularities and capture these understandable patterns. Meanwhile, such a rule-driven perspective also naturally resonates with how humans reason and make decisions. Cognitive studies show that people rarely exhaust all possible options; rather, they often apply simple ”if–else” heuristics to make efficient choices under resource constraints (Gigerenzer et al., 2000; Simon, 1990). For example, if it rains, people carry an umbrella; if the weather is clear, they leave it at home. These everyday heuristics clearly illustrate how humans rely on interpretable rules to balance accuracy and efficiency without overcomplicating decisions. Motivated by this analogy, we design our framework around rule-driven routing, where transparent rules encode observed query-path alignments and are incrementally refined through feedback to balance accuracy, efficiency, and interpretability.
Our framework incorporates three key components: (i) a Rule-Driven Routing Agent, which avoids the inefficiency of static augmentation by evaluating candidate paths with interpretable rules and selecting the most suitable one; (ii) a Rule-Making Expert Agent, which refines these rules using feedback from QA performance, thereby reducing reliance on static, hand-crafted rules and enabling adaptation to different datasets; and (iii) a Path-Level Meta-Cache, which further accelerates inference by reusing routing decisions for repeated or semantically similar queries, eliminating unnecessary agent calls. Together, these components enable adaptive, interpretable, and efficient integration of structured and unstructured knowledge sources. Our contributions are summarized as follows:
  • We conduct analysis of routing between relational databases and document corpora in RAG, highlighting their complementary strengths through motivating experiments.
  • We introduce a rule-driven routing framework that integrates a routing agent, a rule-making expert agent, and a path-level meta-cache, aiming to balance accuracy, interpretability, and efficiency.
  • We evaluate the proposed framework on three QA benchmarks, showing that it achieves consistent improvements over static strategies and alternative routing baselines, while maintaining moderate cost.
@@ 2.1. Hybrid Knowledge Sources for RAG
Retrieval-Augmented Generation (RAG) is an influential framework that enhances the capabilities of large language models (LLMs) by enabling access to external knowledge sources, thereby improving their performance on tasks that require domain-specific information (Borgeaud et al., 2022; Izacard et al., 2023; Lewis et al., 2020). Unlike traditional language models, which generate text based solely on internal parameters, RAG models incorporate a retrieval component that fetches relevant content from external data sources and integrates it into the generation process. Traditional RAG systems primarily utilize large-scale unstructured corpora, such as Wikipedia and web documents, as their external knowledge base (Karpukhin et al., 2020; Ram et al., 2023; Yao et al., 2023). Building on this foundation, more recent work has expanded the spectrum of retrieval sources to further enhance RAG’s capabilities. Some studies integrate knowledge graphs, which provide structured representations of entities and relationships. Knowledge graphs enable more context-aware and entity-centric retrieval, supporting complex information needs and facilitating multi-hop reasoning (Ma et al., 2025; Sarmah et al., 2024; Li et al., 2024). Other approaches incorporate multi-modal data—such as images, videos, and tables—enabling RAG systems to synthesize information across modalities and improving their robustness on heterogeneous, complex queries (Abootorabi et al., 2025; Hu et al., 2023; Chen et al., 2022). In contrast, relational databases store highly structured and precise factual data, supporting efficient updates and fast queries (Harrington, 2016; Codd, 2007). Despite their widespread use in many enterprise and scientific applications, the integration of relational databases as knowledge sources in RAG has received relatively little attention (Seabra et al., 2024). There is a lack of systematic analysis and methodical approaches for combining the complementary strengths of relational databases and unstructured documents within RAG frameworks. Addressing this gap, our work aims to explore and propose strategies for the joint utilization of these two data sources.
@@ 2.2. Routing Mechanisms in RAG
Routing mechanisms have received increasing attention in RAG systems, enabling dynamic and intelligent decision-making about how each query is processed. Unlike fixed pipelines, routing-enhanced RAG architectures employ dedicated modules that analyze the characteristics of incoming queries and adaptively select the most appropriate retrieval and reasoning pathways, thereby improving both efficiency and response quality. A fundamental role of routing in RAG is to determine the appropriate complexity of the retrieval pipeline for each query. While some questions can be resolved through simple, single-hop retrieval, others require more complex, multi-hop reasoning or iterative retrieval procedures (Jeong et al., 2024; Tan et al., 2024; Maekawa et al., 2024). Adaptive routing allows the system to dynamically decide whether external retrieval is needed or if the language model’s parametric knowledge suffices, effectively balancing computational cost and answer accuracy. Another crucial dimension of routing concerns the choice of retrieval methods. Early RAG systems often used simple rule-based or classifier-driven approaches, sending queries with strong lexical overlap to sparse retrievers such as BM25 (Robertson et al., 1994), while directing semantically complex or ambiguous queries to dense retrievers (Karpukhin et al., 2020). More recent advances have introduced lightweight classifiers, neural models, and multi-agent frameworks that can select or combine multiple retrieval strategies, sometimes explicitly modeling the interplay between external retrievals and the language model’s internal knowledge (Şakar and Emekci, 2025; Zhang et al., 2025). Beyond retrieval strategy, routing also governs data source selection. In real-world scenarios, the knowledge necessary to answer a query may reside in document collections, knowledge bases, or other sources. Modern RAG systems increasingly deploy specialized routers—such as domain routers or dynamism routers—to intelligently map queries to the most relevant data sources and to orchestrate hybrid retrieval workflows that capitalize on the complementary advantages of each data type (Wu et al., 2025; Guerraoui et al., 2025; Xia et al., 2025). In this work, we focus on dynamic routing and the integration of structured relational databases with unstructured documents in RAG systems. By recognizing and leveraging the distinct strengths of these sources, we propose a routing mechanism that adaptively selects the knowledge source.
@@ 3. Preliminaries
Database-Enhanced QA. We further consider database-enhanced QA, which leverages structured relational data to provide precise, time-sensitive, and verifiable answers. For a given question $q$, the system first retrieves relevant tables by matching $q$ against a pre-built index of table metadata, including schemas, high-frequency values, and LLM-generated table descriptions. Once candidate tables are identified, factual records $\mathcal{F}_{q}$ are extracted from the database either through advanced text-to-SQL generation methods (Shi et al., 2025; Deng et al., 2022) or via efficient keyword-based filtering. The resulting records serve as immutable evidence and are incorporated into the final prompt: $a=\mathrm{LLM}(q,\mathcal{F}_{q})$. In our implementation, we explicitly instruct the LLM that retrieved database facts are not to be modified or reinterpreted, ensuring that the generated answers remain precise.
@@
Routing Mechanisms. The complementary nature of document and database augmentation presents a new challenge: for each question, how to dynamically select the most effective enhancement pathway to balance answer quality, response latency, and computational cost. We formalize this as a query routing problem: given a question $q$, a routing agent selects an augmentation path $p_{q}\in\mathcal{P}={\texttt{Doc},\texttt{DB},\texttt{Hybrid},\texttt{LLM}}$ according to a policy that optimizes a utility objective combining quality, latency, and cost. The selected pathway determines which information is retrieved and used to construct the prompt for the downstream LLM, which generates the final answer. This design allows the system to flexibly adapt to each query’s needs, leveraging the complementary strengths of both data sources while managing efficiency.
@@ 4.2. Rule-Driven Routing Agent
Expert-Initialized Rules. Our routing design begins with a set of interpretable rules $\mathcal{R}$, initialized by domain experts. Each rule assigns additive scores to candidate augmentation paths based on recognizable query patterns. Formally, we denote by $S_{p}(q)$ the score assigned to path $p$ for query $q$, where higher values indicate stronger preference for selecting that path. A subset of illustrative rules (complete rules are provided in Appendix B) is as follows:
@@ 4.3. Rule-Making Expert Agent
4.4. Path-Level Meta-Cache
While our rule-driven routing agent substantially reduces the inefficiency of naive hybrid augmentation, it still requires an additional agent call for rule evaluation. This extra step introduces some overhead, yet the cost is minor compared to the savings achieved by avoiding redundant hybrid augmentation. Nevertheless, in large-scale deployments, even small latencies can accumulate, motivating further optimization. Since many real-world queries are repeated or semantically similar, caching is a natural choice to further reduce overhead. However, in dynamic scenarios such as frequently updated relational databases, traditional answer-level caches may become unreliable. For example, caching answers to queries like ”What is the revenue growth over the past 7 days?” can easily lead to stale or incorrect results as underlying data change. To overcome this limitation, we propose a path-level meta-cache that operates at the routing decision level rather than the answer level.
(5) $\text{sim}(z_{q_{j}},z_{q^{\prime}})\geq\tau,$
5. Experiment